OpenStudy (anonymous):

Need help with Pythagorean Theorem real quick! How do you find the sum of a triangle when given only C of (A^2 + B^2 = C^2)?

OpenStudy (jojokiw3):

Do you have an angle or anything else?

OpenStudy (jojokiw3):

Or do they tell you what kind of triangle this is?

OpenStudy (anonymous):

Here, I've got a screenshot-- It's kinda hard to explain.

OpenStudy (jojokiw3):

Okay so we're going to assume that's a square, because it looks like one. thus each side is equal to the other.

OpenStudy (jojokiw3):

that means that a^2 = b^2. So then what you can do is find 44^2. Divide that answer by 2. then find the square root of that answer.

OpenStudy (anonymous):

the diagonal of a square is the length of the side times \(\sqrt2\) you have the diagonal, so each side is \(\frac{44}{\sqrt2}\)

OpenStudy (jojokiw3):

There's something I never knew. :P And much better.

OpenStudy (triciaal):